OpenStudy (anonymous):

A 600-kg car is going over a curve with a radius of 120 m that is banked at an angle of 25 degrees with a speed of 30.0 m/s. The coefficient of static friction between the car and the road is 0.30. What is the normal force exerted by the road on the car?

OpenStudy (anonymous):

taking components along x and y direction; mg cos25 - N = mv^2/R and 0.3N=mg sin 25 by subtracting the two equations we get, 1.3 N =mg cos 25+mg sin25-mv^2/R so that should be the value of N.